\medskip

	\begin{thebibliography}{99}
		
		\bibitem{problems} P. J. Cameron and D. B. West, Research problems from the 20th British Combina-
		torial Conference, Discrete Math. 308 (2008), 621--630.

		\bibitem{carlitz}  L. Carlitz, Some cyclotomic matrices, Acta Arith. 5 (1959), 293--308.
		
		\bibitem{chapman} R. Chapman, Determinants of Legendre symbol matrices, Acta Arith. 115 (2004), 231--244.
		\bibitem{evil} R. Chapman, My evil determinant problem, preprint, December 12, 2012, available from http://empslocal.ex.ac.uk/people/staff/rjchapma/etc/evildet.pdf.

\bibitem{ChenSun} D. Chen and Z.-W. Sun, On determinants involving $(\f{j+k}p)\pm(\f{j-k}p)$,
arXiv:2409.08213, 2024.
		
		\bibitem{K1} C. Krattenthaler, Advanced Determinant Calculus. In: Foata D., Han GN. (eds) The Andrews Festschrift. Springer, Berlin, Heidelberg 2001.
		
		\bibitem{K2}  C. Krattenthaler, Advanced determinant calculus: a complement, Linear Algebra
		Appl. 411 (2005), 68--166.

\bibitem{L} D. H. Lehmer, On certain character matrices, Pacific J. Math.  6 (1956), 491--499.
	
		\bibitem{S19} Z.-W. Sun, On some determinants with Legendre symbols entries, Finite Fields Appl. 56 (2019), 285--307.

\bibitem{RJ} Z.-W. Sun,  On some determinants involving the tangent functions, Ramanujan J.
{64} (2024), 309--332.

\bibitem{S24} Z.-W. Sun, Problems and results on determinants involving Legendre symbols, Bull. Math. Soc. Sci. Math. Roumanie, in press. See also arXiv:2405.03626, 2024.
		 
\bibitem{mdl}R. Vrabel, A note on the matrix determinant lemma, Int. J. Pure Appl. Math. 111 (2016), 643--646.
		
\bibitem{V3} M. Vsemirnov, On the evaluation of R. Chapman's ``evil determinant", Linear Algebra Appl. 436 (2012), 4101--4106.
		
\bibitem{V1} M. Vsemirnov, On R. Chapman's ``evil determinant": case $p\equiv1\pmod 4$, Acta Arith. 159 (2013), 331--344.

\bibitem{WW} L.-Y. Wang and H.-L. Wu, On the cyclotomic field $\Q(e^{2\pi i/p})$ and Zhi-Wei Sun's conjecture on $\det M_p$, Finite Fields Appl. 101 (2025), Article 102533.

\bibitem{WWN} L.-Y. Wang, H.-L. Wu and H.-X. Ni,
{\it On a generalization of R. Chapman's ``evil determinant"}, preprint, Rev. Real Acad. Cienc. Exactas Fis. Nat. Ser. A-Mat. 120 (2026), Article 43.
	
	\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage{enumerate, latexsym, amsmath, amsfonts, amssymb, amsthm, graphicx, color}
\usepackage{hyperref}
\hypersetup{hypertex=true,
	colorlinks=true,
	linkcolor=blue,
	anchorcolor=blue,
	citecolor=blue}
\usepackage{amsmath}
\def\pmod #1{\ ({\rm{mod}}\ #1)}
\def\i{\mathrm{i}}

\def\Q{\Bbb Q}

\def\l{\left}
\def\r{\right}
\def\bg{\bigg}
\def\({\bg(}
\def\){\bg)}
\def\t{\text}
\def\f{\frac}

\def\adj{{\rm adj}}

\def\ls{\le}

\def\ve{\varepsilon}

\def\eq{\equiv}

\def\u{{\bf u}}
\def\v{{\bf v}}
\def\Proof{\noindent{\it Proof}}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}

\newtheorem{lemma}{Lemma}

\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{remark}{Remark}

\makeatletter
\@namedef{subjclassname@2010}{%
	\textup{2010} Mathematics Subject Classification}
\makeatother
\vspace{4mm}

\allowdisplaybreaks[1]
\begin{document}	
	\hbox{Preprint}
\medskip

	\title[A determinant involving Legendre symbols]
{On a determinant involving linear combinations of Legendre symbols}
	
	\author[K. Liu, Z.-W. Sun and L.-Y. Wang]{Keqin Liu*, Zhi-Wei Sun and Li-Yuan Wang}
	
	\address {(Keqin Liu) School of Mathematics and Physics, Xi'an Jiaotong-Liverpool University,
	Suzhou, 215123, People's Republic of China} \email{Keqin.Liu@xjtlu.edu.cn}

\address {(Zhi-Wei Sun) School of Mathematics, Nanjing University\\
	Nanjing, 210093, People's Republic of China} \email{zwsun@nju.edu.cn}

\address {(Li-Yuan Wang) School of Physical and Mathematical Sciences, Nanjing Tech University, Nanjing, 211816, People's Republic of China}
	\email{\tt lywang@njtech.edu.cn}

	\begin{abstract}
   In this paper, we prove a conjecture of the second author by evaluating the determinant
  $$\det\left[x+\left(\frac{i-j}p\right)+\left(\frac ip\right)y+\left(\frac jp\right)z+\left(\frac{ij}p\right)w\right]_{0\ls i,j\ls(p-3)/2}$$
  for any odd prime $p$, where $(\frac{\cdot}p)$ denotes the Legendre symbol.
  In particular, the determinant is equal to $x$ when $p\equiv 3\pmod4$. 
	\end{abstract}
	
	\thanks{2020 {\it Mathematics Subject Classification}.
		Primary 11A15, 11C20; Secondary 15A15.
		\newline\indent {\it Keywords}. Determinant, Legendre symbol, quadratic field.
		\newline \indent
		The work of Keqin Liu was supported by Leadership Talent Program (Science and Education) of SIP (Grant No. KJQ2024202). The work of Zhi-Wei Sun was supported by the National Natural Science Foundation of China (Grant Nos. 12371004, 12201291) and the Natural Science Foundation of the Higher Education Institutions of Jiangsu Province (21KJB110001). The work of Li-Yuan Wang was supported by the National Natural Science Foundation of China (Grant No. 12201291).
		\newline\indent{*Corresponding author.}
        }
	\maketitle
	\section{Introduction}	
	\setcounter{lemma}{0}
	\setcounter{theorem}{0}
	\setcounter{corollary}{0}
	\setcounter{remark}{0}
	\setcounter{equation}{0}
	\setcounter{conjecture}{0}

For a matrix $A=[a_{ij}]_{1\ls i,j\ls n}$ over a field, we use $\det A$ or $|a_{ij}|_{1\ls i,j\ls n}$
to denote its determinant. If $a_{ji}=-a_{ij}$ for all $i,j=1,\ldots,n$, then we say that $A$ is skew-symmetric, and in this case we have
$$\det A=|a_{ji}|_{1\ls i,j\ls n}=|-a_{ij}|_{1\ls i,j\ls n}=(-1)^n\det A.$$
In particular, $\det A=0$ if $A$ is skew-symmetric and $n$ is odd.

Let $p$ be an odd prime, and let $(\f{\cdot}p)$ be the Legendre symbol.
For any integer $m$, Lehmer \cite{L} and Carlitz \cite{carlitz} found all the eigenvalues of the matrices
$$\l[x+\l(\f{j+k+m}p\r)\r]_{1\ls j,k\ls p-1}\ \l[x+\l(\f{j-k+m}p\r)\r]_{1\ls j,k\ls p-1}$$
respectively.

Let $\ve_p$ and $h_p$ denote the fundamental unit and class number
	of the real quadratic field $\mathbb{Q}(\sqrt{p})$, respectively.
For convenience, we write
\begin{align}\label{abp}
	 \varepsilon_p^{h_p}=a_p+b_p\sqrt{p}\ \ \t{with}\ a_p,b_p\in\mathbb{Q},
\end{align}
and
\begin{align}\label{ab'p}
	 \varepsilon_p^{(2-(\frac{2}{p}))h_p}=a_p'+b_p'\sqrt{p}\ \ \t{with}\ a_p',b_p'\in\mathbb{Q}.
\end{align}
In 2004, motivated by coding theory, Chapman \cite{chapman} proved that
 \begin{align*}&\ \l|x+\l(\f{i+j-1}p\r)\r|_{1\ls i,j\ls(p+1)/2}
 \\=&\ \begin{cases}(\f 2p)2^{(p-1)/2}(pb_px-a_p)&\t{if}\ p\eq1\pmod4,
 \\2^{(p-1)/2}&\t{if}\ p>3\ \t{and}\ p\eq3\pmod4.\end{cases}
 \end{align*}
 As
 $$\f{p+1}2-i+\f{p+1}2-j-1=p-i-j\eq-(i+j)\pmod p,$$
 we see that
 $$\l|\l(\f{-1}p\r)x+\l(\f{i+j}p\r)\r|_{0\ls i,j\ls(p-1)/2}=\l|x+\l(\f{i+j-1}p\r)\r|_{1\ls i,j\ls(p+1)/2}.$$
 Chapman's ``evil determinant conjecture" (cf. \cite{problems,evil}) states that
	\begin{align}\label{evil}
		 \left|\left(\frac{i-j}{p}\right)\right|_{0\le i,j\le (p-1)/2}=\begin{cases}
		 	-a_p' & \mbox{if}\ p\equiv 1\pmod4,\\
		 	1    & \mbox{if}\ p\equiv 3\pmod4,
		 \end{cases}
	\end{align}
 this was  confirmed by Vsemirnov \cite{V3,V1}  via matrix decomposition.

Sun \cite{S19} defined $M_p$ as the matrix obtaining from $[(\f{i-j}p)]_{0\ls i,j\ls (p-1)/2}$ via replacing all the entries in the first row by $1$, and conjectured that
\begin{equation}\label{Mp}\det M_p=\begin{cases}(-1)^{(p-1)/4}&\text{if}\ p\equiv 1\pmod4,
\\(-1)^{(h(-p)-1)/2}&\text{if}\ p>3\ \t{and}\ p\equiv3\pmod4,\end{cases}
\end{equation}
where $h(-p)$ denotes the class number of the imaginary quadratic field $\mathbb Q(\sqrt{-p})$
which is known to be odd. This conjecture was recently confirmed by Wang and Wu \cite{WW}.

Recently, Sun \cite{S24} determined
$$\l|x+\l(\f{i\pm j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z\r|_{0\ls i,j\ls(p-1)/2},$$
and posed many conjectures on determinants involving linear combinations of Legendre symbols.

In this paper, we establish the following result conjectured by Sun \cite[Conjecture 3.2(ii)]{S24}.

	\begin{theorem}\label{Th1} Let $p$ be an odd prime, and let $a_p'$ and $b_p'$ be
given by  $(\ref{abp})$.

{\rm (i)} In the case $p\eq1\pmod 4$, we have
\begin{equation}\label{main1}
 \begin{aligned}&\ \l|x+\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z+\l(\f{ij}p\r)w\r|_{0\ls i,j\ls(p-3)/2}
\\&\qquad\ =((y+1)(z+1)-wx)\l(\f 2p\r)b_p'-a_p'x.
\end{aligned}
\end{equation}

{\rm (ii)} Provided that $p\eq3\pmod4$, we have
\begin{equation}\label{main3} \l|x+\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z+\l(\f{ij}p\r)w\r|_{0\ls i,j\ls(p-3)/2}=x.
\end{equation}
\end{theorem}

It is easy to verify \eqref{main3} for $p=3$. From now on, we let $p$ be a fixed prime greater than $3$.

In the next section, we will reduce the two parts of Theorem \ref{Th1} to
 the identity
\begin{equation}\label{simple}\l|x+\l(\f{i-j}p\r)\r|_{1\ls i,j\ls(p-1)/2}=\begin{cases}
(\frac{2}{p})b'_p-a_p'x&\t{if}\ p\eq1\pmod4,
\\x&\t{if}\ p\eq3\pmod4.
\end{cases}
\end{equation}
This identity in the case $p\eq3\pmod4$ first appeared as a conjecture
in \cite{S19}. In Section 3, we will prove \eqref{simple} in the case $p\eq3\pmod4$.
Based on some lemmas in Section 4, we are going to prove \eqref{simple} for the case $p\eq1\pmod4$
in Section 5.

Throughout this paper, for a matrix $A$ we use $A^T$ to denote the transpose of $A$.
For a matrix $A=[a_{ij}]_{1\ls i,j\ls n}$ over a field, its adjugate matrix is given by $\adj(A)=[A_{ji}]_{1\ls i,j\ls n}$, where $A_{ji}$ is the cofactor of the entry $a_{ji}$ in $A$.

	\section{Reduction of Theorem \ref{Th1} to the identity \eqref{simple}}
	\setcounter{lemma}{0}
	\setcounter{theorem}{0}
	\setcounter{corollary}{0}
	\setcounter{remark}{0}
	\setcounter{equation}{0}
	\setcounter{conjecture}{0}

The following basic lemma can be found in \cite[Lemma 2.1]{RJ}.

\begin{lemma} \label{Lem2.1} Let $A=[a_{ij}]_{0\ls i,j\ls m}$ be a matrix over a field. Then
\begin{equation}\label{ab}\det[x+a_{ij}]_{0\ls i,j\ls m}-\det[a_{ij}]_{0\ls i,j\ls m}
=x\det[b_{ij}]_{1\ls i,j\ls m},
\end{equation}
where $b_{ij}=a_{ij}-a_{i0}-a_{0j}+a_{00}$.
\end{lemma}

Fix an odd prime $p$. Applying Lemma \ref{Lem2.1}, we obtain
\begin{equation}\label{linear}\begin{aligned}&\ \l|x+\l(\f{i-j}p\r)\r|_{1\ls i,j\ls(p-1)/2}=\l|x+\l(\f{i-j}p\r)\r|_{0\ls i,j\ls(p-3)/2}
\\=&\ \l|\l(\f{i-j}p\r)\r|_{0\ls i,j\ls(p-3)/2}
+x\l|\l(\f{i-j}p\r)-\l(\f ip\r)-\l(\f{-j}p\r)\r|_{1\ls i,j\ls(p-3)/2}.
\end{aligned}\end{equation}
On the other hand, by \cite[Theorem 1.2(ii)]{S24} we have
\begin{align*}&\ \l(\f{-1}p\r)\l|x+\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z\r|_{0\ls i,j\ls(p-3)/2}
\\=&\ (y+1)\l(z+\l(\f{-1}p\r)\r)\l|\l(\f{i-j}p\r)\r|_{0\ls i,j\ls(p-3)/2}
\\&\ +\l(\f{-1}p\r)x\l|\l(\f{i-j}p\r)-\l(\f ip\r)-\l(\f{-j}p\r)\r|_{1\ls i,j,\ls(p-3)/2}.
\end{align*}
Thus, from \eqref{simple} we can deduce the more general conclusion:
\begin{equation}\label{-w}
\begin{aligned}&\ \l|x+\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z\r|_{0\ls i,j\ls(p-3)/2}
\\=&\ \begin{cases}(y+1)(z+1)(\f 2p)b_p'-a_p'x&\t{if}\ p\eq1\pmod4,
\\x&\t{if}\ p\eq3\pmod4.
\end{cases}
\end{aligned}
\end{equation}
Combining this with \cite[Theorem 1.2(i)]{S24}, we see that in the case $p\eq1\pmod4$
the equality \eqref{simple} implies that
\begin{equation}\label{no-w}
\begin{aligned}&\ \l|\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z+\l(\f{ij}p\r)w\r|_{0\ls i,j\ls(p-3)/2}
\\=&\ \l|\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z\r|_{0\ls i,j\ls(p-3)/2}
\\=&\ (y+1)(z+1)\l(\f 2p\r)b_p'.
\end{aligned}
\end{equation}

Now we consider the case $p\eq1\pmod4$. Define
$$A(x,y,z,w)=\l[x+\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z+\l(\f{ij}p\r)w\r]_{0\ls i,j\ls(p-3)/2}.$$
If \eqref{simple} holds, then
\begin{gather*}|A(0,0,0,0)|=\l(\f 2p\r)b_p'\not=0,\ |A(1,0,0,0)|=\l(\f 2p\r)b_p'-a_p',
\\ |A(0,1,0,0)|=|A(0,0,1,0)|=2\l(\f 2p\r)b_p'
\ \t{and}\ |A(0,0,0,1)|=\l(\f 2p\r)b_p'
\end{gather*}
with the aids of \eqref{-w} and \eqref{no-w}, hence by applying \cite[Theorem 3.1]{ChenSun}
we obtain
$$|A(x,y,z,w)|=((y+1)(z+1)-wx)\l(\f 2p\r)b_p'-a_p'x.$$

For $i,j=0,\ldots,(p-3)/2$, let
$$a_{ij}=\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z+\l(\f{ij}p\r)w.$$
By \cite[Theorem 2.1(i)]{S24}, $\l|a_{ij}\r|_{0\ls i,j\ls (p-3)/2}$
does not depend on $w$.
Combining this with Lemma \ref{Lem2.1}, we get
\begin{align*} &\ \l|x+a_{ij}\r|_{0\ls i,j\ls (p-3)/2}
\\ =&\ \l|\l(\f{i-j}p\r)+\l(\f ip\r)y+\l(\f jp\r)z\r|_{0\ls i,j\ls (p-3)/2}+x|b_{ij}|_{1\ls i,j\ls(p-3)/2},
\end{align*}
where
\begin{align*}b_{ij}&=a_{ij}-a_{i0}-a_{0j}+a_{00}
\\&=a_{ij}-\l(\f ip\r)(y+1)-\l(\f jp\r)\l(z+\l(\f{-1}p\r)\r)+0
\\&=\l(\f{i-j}p\r)-\l(\f ip\r)-\l(\f{-j}p\r)+\l(\f{ij}p\r)w.
\end{align*}
In light of Lemma \ref{Lem2.1},
\begin{align*}|b_{ij}|_{1\ls i,j\ls(p-3)/2}&=\l|\l(\f{ij(i-j)}p\r)-\l(\f jp\r)-\l(\f{-i}p\r)+w\r|_{1\ls i,j\ls(p-3)/2}
\\&=\l|\l(\f{ij(i-j)}p\r)-\l(\f jp\r)-\l(\f{-i}p\r)\r|_{1\ls i,j\ls(p-3)/2}
\\&\quad +w|c_{ij}|_{2\ls i,j\ls(p-3)/2},
\end{align*}
where
\begin{align*}c_{ij}=&\ \l(\f{ij(i-j)}p\r)-\l(\f jp\r)-\l(\f{-i}p\r)
\\&\ -\l(\f{j(1-j)}p\r)+\l(\f jp\r)+\l(\f{-1}p\r)
\\&\ -\l(\f{i(i-1)}p\r)+1+\l(\f{-i}p\r)-\l(\f1p\r)-\l(\f{-1}p\r)
\\=&\ \l(\f{ij(i-j)}p\r)-\l(\f{i(i-1)}p\r)-\l(\f{j(1-j)}p\r).
\end{align*}

Now assume that $p\eq3\pmod4$. Then $|c_{ij}|_{2\ls i,j\ls(p-3)/2}$ vanishes since $[c_{ij}]_{2\ls i,j\ls(p-3)/2}$ is a skew-symmetric matrix of odd order.
Hence $|b_{ij}|_{1\ls i,j\ls(p-3)/2}$ does not depend on $w$.
Therefore
$|x+a_{ij}|_{0\ls i,j\ls(p-3)/2}$ does not depend on $w$, and its value is $x$ provided  \eqref{-w}.

In view of the above, we have reduced Theorem \ref{Th1} to the identity \eqref{simple}.

	\section{Proof of \eqref{simple} in the case $p\eq3\pmod4$}
	\setcounter{lemma}{0}
	\setcounter{theorem}{0}
	\setcounter{corollary}{0}
	\setcounter{remark}{0}
	\setcounter{equation}{0}
	\setcounter{conjecture}{0}

Let $p>3$ be a prime with $p\eq3\pmod4$.
As $[(\f{i-j}p)]_{0\ls i,j\ls(p-3)/2}$
is a skew-symmetric matrix of odd order, we have
\begin{equation}\label{zero}\l|\l(\f{i-j}p\r)\r|_{1\ls i,j\ls(p-1)/2}=\l|\l(\f{i-j}p\r)\r|_{0\ls i,j\ls(p-3)/2}=0.
\end{equation} Combining this with \eqref{linear}, we see that
$$\l|x+\l(\f{i-j}p\r)\r|_{1\ls i,j\ls (p-1)/2}=mx$$
for an integer $m$ not depending on $x$. Thus
$$\l|x+\l(\f{i-j}p\r)\r|_{1\ls i,j\ls (p-1)/2}=x$$
if
\begin{equation}
\label{1+}\l|1+\l(\f{j-i}p\r)\r|_{1\ls i,j\ls (p-1)/2}=1.
\end{equation}
So it suffices to prove \eqref{1+}.

Let $M_p=[m_{ij}]_{0\ls i,j\ls (p-1)/2}$ be the matrix obtained from $[(\f{i-j}p)]_{0\ls i,j\ls(p-1)/2}$
 via replacing all the entries in the first row by $1$, and let $M_p^*=[m_{ij}^*]_{0\ls i,j\ls(p-1)/2}$ be the matrix $[m_{(p-1)/2-i,(p-1)/2-j}]_{0\ls i,j\ls (p-1)/2}$. Then
$$\det M_p^*=\det M_p=(-1)^{(h(-p)-1)/2}$$
by \eqref{Mp}.
For each $j=0,\ldots,(p-1)/2$, we have $$m_{ij}^*=\begin{cases}(\f{j-i}p)&\t{if}\ 0\ls i<(p-1)/2,\\1&\t{if}\ i=(p-1)/2.\end{cases}$$

As $|(\f{j-i}p)|_{0\ls i,j\ls (p-1)/2}=1\not=0$, by Cramer's rule there are rational numbers
$c_0,\ldots,c_{(p-1)/2}$ such that
$$\sum_{i=0}^{(p-1)/2}\l(\f{j-i}p\r)c_i=1\quad\t{for all}\ j=0,\ldots,\f{p-1}2.$$
Define a lower triangular matrix $A=[a_{ij}]_{0\ls i,j\ls(p-1)/2}$ by
$$a_{ij}=\begin{cases}1&\t{if}\ i=j\in\{0,\ldots,(p-3)/2\},
\\c_j&\t{if}\ i=(p-1)/2\ \t{and}\ 0\ls j\ls(p-1)/2,
\\0&\t{otherwise}.\end{cases}$$
Then
$$A\l[\l(\f{j-i}p\r)\r]_{0\ls i,j\ls(p-1)/2}=M_p^*$$
and hence
\begin{equation}
\label{A}\det A=\det A\cdot\l|\l(\f{j-i}p\r)\r|_{0\ls i,j\ls(p-1)/2}=\det M_p^*=(-1)^{(h(-p)-1)/2}.
\end{equation}

Observe that  $M_p^*A^T$ coincides with the matrix $B=[b_{ij}]_{0\ls i,j\ls(p-1)/2}$ with
$$b_{ij}=\begin{cases}(\f{j-i}p)&\t{if}\ i,j\in\{0,\ldots,(p-3)/2\},
\\-1&\t{if}\ i<j=(p-1)/2,
\\1&\t{if}\ i=(p-1)/2>j,\\c&\t{if}\ i=j=(p-1)/2,
\end{cases}$$
where $c=\sum_{i=0}^{(p-1)/2}c_i$. Therefore
\begin{equation}\label{B}\det B=\det M_p^* \cdot \det A=1
\end{equation}
by \eqref{A}.
Let $B_*$ be the matrix obtained from $B$ by replacing the right-bottom entry $c$ by $1$.
Then $\det B_*=\det B=1$ since $|(\f{j-i}p)|_{0\ls i,j\ls(p-3)/2}=0$ by \eqref{zero}.
Via adding the last row of $B_*$ to all previous rows, we see that
$$\det B=\l|1+\l(\f{j-i}p\r)\r|_{0\ls i,j\ls(p-3)/2}=\l|1+\l(\f{j-i}p\r)\r|_{1\ls i,j\ls(p-1)/2}.$$
Combining this with \eqref{B}, we immediately get the desired identity \eqref{1+}.

In view of the above, we have completed our proof of \eqref{simple} in the case $p\eq3\pmod4$.

	\section{Some lemmas}
	\setcounter{lemma}{0}
	\setcounter{theorem}{0}
	\setcounter{corollary}{0}
	\setcounter{remark}{0}
	\setcounter{equation}{0}
	\setcounter{conjecture}{0}

To prove \eqref{simple} for primes $p\eq1\pmod4$, we need
	several lemmas.

	\begin{lemma}[The Matrix-Determinant Lemma]\label{mdl}
		Let $H$ be an $m\times m$ matrix over a field $F$, and let $\u,\v$ be two $m$-dimensional column vectors with entries in $F$. Then we have
		\begin{equation*}
			\det (H+\u\v^T)=\det H+ \v^T \adj(H)\u.
		\end{equation*}		
	\end{lemma}
	
\begin{remark} This lemma is well-known. One may consult \cite{mdl} for a  proof.
\end{remark}

The following known lemma can be found in \cite[Theorem 3]{V1}.

\begin{lemma} \label{uv} We have
	\begin{align*}
		&\	\det \bigg[\frac{x_i+y_j}{1+x_iy_j}\bigg]_{1\le i,j\le m}  \\
		= &\ \frac{1}{2}\(\prod_{i=1}^{m}(  1+x_i )(  1+y_i )+(-1)^m \prod_{i=1}^{m}(  1-x_i )(  1-y_i )\)  \\
		&\ \cdot\prod_{1\le i<j\le m}^{}(x_i-x_j)(y_j-y_i)\cdot\prod_{i=1}^{m}  \prod_{j=1}^{m}(1+x_iy_j)^{-1}.
	\end{align*}
\end{lemma}

\begin{lemma}\label{a+-b} Let $p$ be a prime with $p\eq1\pmod4$, and set $\zeta =e^{2\pi \i/p}$.  Then
	\begin{equation}
	  \prod_{j=1}^{n}\( 1+ \(\frac{j}{p}\)\zeta^{-j} \)^2 =(-1)^{n/2}\zeta^{-n(n+1)/2}(b_p' p+ a_p' \sqrt{p}),
\end{equation}
and
\begin{equation}
 \prod_{j=1}^{n}\( 1- \(\frac{j}{p}\)\zeta^{-j} \)^2 =(-1)^{n/2}\zeta^{-n(n+1)/2}(b_p' p- a_p' \sqrt{p}).
	\end{equation}
\end{lemma}
\Proof. Clearly, the desired results follow from the following two  identities:
	\begin{align}\label{bp}
	\frac{1}{2} \(   \prod_{j=1}^{n}\( 1+ \(\frac{j}{p}\)\zeta^{j} \)^2+\prod_{j=1}^{n}\( 1- \(\frac{j}{p}\)\zeta^{j} \)^2 \)=(-1)^{n/2}\zeta^{n(n+1)/2} b_p' p,
\end{align}
\begin{align}\label{ap}
	\frac{1}{2} \(   \prod_{j=1}^{n}\( 1+ \(\frac{j}{p}\)\zeta^{j} \)^2-\prod_{j=1}^{n}\( 1- \(\frac{j}{p}\)\zeta^{j} \)^2 \)=(-1)^{n/2}\zeta^{n(n+1)/2} a_p' \sqrt{p}.
\end{align}
 Actually, (\ref{ap}) can be found in  \cite[Lemma 2]{V1},  and (\ref{bp})  can be proved similarly. \qed

	For convenience, we introduce Vsemirnov's notations which will be used soon.

Let $p$ be a prime with $p\eq1\pmod4$, and set
$\zeta =e^{2\pi i/p}$.
As in \cite{V1}, we define matrices $D,U, V$ of order $(p+1)/2$ whose
$(i,j)$ entries $(0\ls i,j\ls (p-1)/2)$ are as follows:
$$	d_{ij}=\begin{cases}\prod_{0\le k\le n \atop k\ne i}^{}\frac{1}{\zeta^{2i}-\zeta^{2k}}, &\t{if}\ i=j,\\0&\t{otherwise},\end{cases}$$
$$
	u_{ij}=\frac{(\frac{i}{p})\zeta^{-j-2i}+(\frac{j}{p})\zeta^{-2j-i}}  { \zeta^{-i-j}+(\frac{i}{p})(\frac{j}{p}) }\ \ \t{and}\ \
	v_{ij}=\zeta^{2ij}.$$
Vsemirnov \cite[Theorem 2]{{V1}} proved that
$$\left|\left(\frac{j-i}{p}\right)\right|_{0\le i,j\le \frac{p-1}{2}}=\lambda VDUDV,$$	
where $\lambda=(\frac{2}{p})\sqrt{p}\,\zeta^{(p-1)/4}$.

\begin{lemma}\label{New-decompose} Let $p$ be a prime with $p\eq1\pmod4$. For the matrix
$$C(x):=\left[x+\left(\frac{j-i}{p}\right)\right]_{0\le i,j\le \frac{p-1}{2}},
$$
	we have
	\begin{align}\label{cx}
C(x) =\lambda VD\tilde{ U}DV,
	\end{align}
where the left top entry of $ \tilde{U} $is $ (\frac{2}{p}) \sqrt{p}\, x $ and all the other entries are the same as $U$.
\end{lemma}
\Proof.
 It suffices to show that
 \begin{equation}\label{x11}
 	 \lambda VD(\tilde{ U}-U)DV   	=x  ( 1  ,  1,1,\ldots,1  )^T ( 1  ,  1,1,\ldots,1  ).
 \end{equation}
Note that $ \tilde{U}-U= (\frac{2}{p}) \sqrt{p}  x  ( 1  ,  0,0,\ldots,0  )^T  ( 1  ,  0,0,\ldots,0  )$
and
 \begin{align}
 	 ( 1  ,  0,0,\ldots,0  ) DV =d_{00} ( 1  ,  1,1,\ldots,1  ).
 \end{align}
By symmetry,
 \begin{align}
VD  ( 1  ,  0,0,\ldots,0  )^T =d_{00}  ( 1  ,  1,1,\ldots,1  )^T.
\end{align}

Set $n=(p-1)/2$. Then
\begin{align*}
	\frac{1}{	d_{00}^{2} }&= \prod_{0< k\le n }^{}  (1-\zeta^{2k})^2
	 \\
	&=  \prod_{k=1 }^{n}  (-\zeta^{2k})  (1-\zeta^{-2k})  (1-\zeta^{2k})     \\
	&=  (-1)^n\zeta^{n(n+1)}  \prod_{k=1 }^{n}  (1-\zeta^{p-2k}) (1-\zeta^{2k})
\\&=  \zeta^{(p+1)(p-1)/4}   \prod_{r=1 }^{p-1}  (1-\zeta^{r})   \\&= \zeta^{(p-1)/4}\lim_{x\to1}\f{x^p-1}{x-1}=p\zeta^{(p-1)/4}.
\end{align*}
Thus
$d_{00}^{2}=p^{-1} \zeta^{-(p-1)/4}.$
Combining this with the first paragraph in this proof, we obtain the desired result.
 \qed

	\section{Proof of \eqref{simple} in the case $p\eq1\pmod4$}
\setcounter{lemma}{0}
\setcounter{theorem}{0}
\setcounter{corollary}{0}
\setcounter{remark}{0}
\setcounter{equation}{0}
\setcounter{conjecture}{0}

Let  $p$ be a prime with $p\equiv 1\pmod4$, and set $n=(p-1)/2$.
Define
	$$R(x):=\left[x+\left(\frac{j-i}{p}\right)\right]_{1\le i,j\le n}.
$$
Since  $\det R(x)$ is the cofactor of the left top entry of $C(x)$, we have
$$\det R(x)= ( 1  ,  0,0,\ldots,0  ) \adj (C(x))  ( 1  ,  0,0,\ldots,0  )^T.$$
By Lemma \ref{New-decompose},
\begin{align}
	  \adj (C(x))=\lambda^n\cdot	\adj (V) \cdot	\adj(D) \cdot\adj (\tilde{ U})  \cdot 	\adj (D)\cdot	 \adj (V) . \nonumber
\end{align}
So
\begin{align}
	 \det R(x)=&\lambda^n\cdot ( 1  ,  0,0,\ldots,0  )\cdot 	\adj (V) \cdot \adj (D) \cdot\adj (\tilde{ U})    \nonumber \\
	 &	\cdot	\adj (D)\cdot	\adj (V)  \cdot  ( 1  ,  0,0,\ldots,0  )^T.
\end{align}

 Recall that $\adj (V)=[V_{ji}]_{0\le i,j\le m}$ with
 $V_{ji}$ the cofactor of $v_{ji}$ in the matrix $V$.
It is easy to see that
\begin{align}
	 ( 1  ,  0,0,\ldots,0  )\cdot 	\adj (V) \cdot	\adj (D)=	\det (D) \( \frac{ V_{00}}{d_{00}} ,   \frac{ V_{10}}{d_{11}}, \frac{ V_{20}}{d_{22}},\ldots, \frac{ V_{n0}}{d_{nn}} \).
\end{align}

Let $0\le k\le n $.  As determinants of Vandermonde's type can be evaluated, we get
	\begin{align}
	V_{k0}&=V_{0k}=(-1)^{0+k}\cdot\det [\zeta^{2ij}]_{1\le i\le n\atop  0\le j\le n, j\ne k}  \nonumber    \\
		&=(-1)^{k}\cdot     \prod_{j=0\atop j\ne k}^{n} \zeta^{2j}  \cdot   \det  [\zeta^{2(i-1)j}]_{1\le i\le n\atop  0\le j\le n, j\ne k}   \nonumber \\
	&= \zeta^{n(n+1)}  \cdot  \zeta^{-2k} \cdot \frac{\prod_{0\le i<j \le n}(\zeta^{2j}-\zeta^{2i})}{\prod_{0\le j\le n\atop j\ne k}(\zeta^{2k}-\zeta^{2j})}. \nonumber
\end{align}
Since
 $$d_{kk} \prod_{0\le j\le n\atop j\ne k}(\zeta^{2k}-\zeta^{2j})=1
 \ \t{and}\ \det V=\prod_{0\le i<j \le n}(\zeta^{2j}-\zeta^{2i}),$$
 we have
 \begin{align}
  \frac{ V_{k0}}{d_{kk}}=   \zeta^{n(n+1)}\cdot \det V\cdot \zeta^{-2k}.\nonumber
 	\end{align}

 Define
 \begin{equation}
 	\boldsymbol{\alpha}:=    ( 1  , \zeta^{-2}, \zeta^{-4},\ldots,\zeta^{-2n} )^T.\nonumber
 \end{equation}
 Then
 $$( 1  ,  0,0,\ldots,0  )\cdot 	\adj (V) \cdot	\adj(D)=	\zeta^{n(n+1)}\cdot \det D\cdot  \det V	 \cdot \boldsymbol{\alpha}^T.$$
 By symmetry,
 $$\adj (D)\cdot	\adj (V) 	( 1  ,  0,0,\ldots,0  )^T=	\zeta^{n(n+1)}\cdot \det D\cdot  \det V\cdot 	 \boldsymbol{\alpha}.$$
Hence
\begin{equation}\label{Rx}
	\det R(x)= p^{n/2}   \zeta^{\frac{n^2}{2 }  +2n(n+1)}\cdot  (\det V)^2 \cdot (\det D)^2\cdot  \boldsymbol{\alpha}^T \cdot   	\adj (\tilde{ U})  \cdot\boldsymbol{\alpha}.
\end{equation}

  Now it remains  to evaluate $\boldsymbol{\alpha}^T \cdot   	\adj (\tilde{ U})  \cdot\boldsymbol{\alpha}$.  By Lemma \ref{mdl}, we have
  \begin{equation}\label{alphaUalpha}
  	\boldsymbol{\alpha}^T \cdot   	\adj (\tilde{ U})\cdot\boldsymbol{\alpha}=\det (\boldsymbol{\alpha} \boldsymbol{\alpha}^T +\tilde{ U}	)- \det \tilde{ U} .
  \end{equation}
Set $x_i=(\frac{i}{p})\zeta^{-i}$ for $1\le i\le n$. Then
   \begin{equation*}
  	\boldsymbol{\alpha}^T=  ( 1  , x_1^{2}, x_2^{2},\ldots,x_n^{2} ).
  \end{equation*}
 Define
\begin{displaymath}
	G =	\left[ \begin{array}{cccccccc}
		1& & \  &  \\
		&	 x_1^{-1}&    &  \\
		& &	 x_2^{-1}&    &  \\
		&	& & \ddots & \\
		&	& &  &	 x_n^{-1}\\
	\end{array} \right],
\end{displaymath}
and write $y=(\frac{2}{p}) \sqrt{p} x $.
Then
\begin{equation}\label{u}
	G  \tilde{ U}	 G =\left[ \begin{array}{c|cccc}
		y & 1& \cdots & 1 \\
		\hline
		1	 &    & \\
		\vdots & &\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n} \\
		1 &  & \\
	\end{array} \right].
\end{equation}

Let's  turn to simplify $\det (\boldsymbol{\alpha} \boldsymbol{\alpha}^T +\tilde{ U}	)$.
It is easy to verify that
\begin{align}
&G^2 (\boldsymbol{\alpha} \boldsymbol{\alpha}^T +\tilde{ U}	) G^2  	 \nonumber\\
=	&
\left[ \begin{array}{c|cccc}
		y+1& 1+x_1^{-1}  &  1+x_2^{-1}   & \cdots &  1+x_n^{-1}  \nonumber  \\
		\hline
	 1+x_1^{-1}   & & & \\
	 1+x_2^{-1}  	 & &   & \\
		\vdots & && \bigg[1+\frac{x_i^{-1}+x_j^{-1}}{1+x_ix_j}\bigg]_{1\le i,j\le n} &  \nonumber\\
	 1+x_n^{-1}  & & & \\
	\end{array} \right].
\end{align}
Hence
\begin{align}
	&\ \det (	G^2 (\boldsymbol{\alpha}  \boldsymbol{\alpha}^T +\tilde{ U}	) G^2 )    \nonumber \\
	=&\	\det\left[\begin{array}{cc|cccc}
			1 & 1 &1&1& \cdots & 1  \nonumber \\
		0& y+1  &  1+x_1^{-1}   & 1+x_2^{-1}   & \cdots &  1+x_n^{-1}   \nonumber \\
		\hline
	0& 	1+x_1^{-1}   & & & \\
	0& 	1+x_2^{-1}  	 & &   & \\
\vdots & 	\vdots & & &\bigg[1+\frac{x_i^{-1}+x_j^{-1}}{1+x_ix_j}\bigg]_{1\le i,j\le n}    \nonumber \\
	0& 	1+x_n^{-1}  & & &     \nonumber\\
	\end{array} \right]     \nonumber\\
=&\	\det L,  \nonumber
\end{align}
where
\begin{align}
	 L=\left[  \begin{array}{cc|cccc}
	 	1 & 1 &1& 1&\cdots & 1     \nonumber \\
	 	-1& y  &  x_1^{-1}   &  x_2^{-1}   &\cdots &  x_n^{-1}     \nonumber \\
	 	\hline
	 	-1& 	x_1^{-1}   & &   \nonumber \\
	 	-1& 	x_2^{-1}  	 &    &     \nonumber\\
	 	\vdots & 	\vdots & &&\bigg[\frac{x_i^{-1}+x_j^{-1}}{1+x_ix_j}\bigg]_{1\le i,j\le n}    \nonumber \\
	 	-1& 	x_n^{-1}  &  & \\
	 \end{array} \right].
\end{align}

Define the diagonal matrix
\begin{displaymath}
	\tilde{G }=	\left[ \begin{array}{cccccccc}
		1& & \  &  \\
			&1& & \  &  \\
	&	&	 x_1^{-1}&    &  \\
	&	& &	 x_2^{-1}&    &  \\
	&	&	& & \ddots & \\
	&	&	& &  &	 x_n^{-1}\\
	\end{array} \right].
\end{displaymath}
Then
  \begin{align}
  	\tilde{G }^{-1}L\tilde{G }^{-1}=	\left[\begin{array}{cc|cccc}
  		1 & 1 & x_1&x_2& \cdots & x_n \\
  		-1& y  &1  &1 & \cdots &  1 \\
  		\hline
  		-x_1& 1  & & \\
  		-x_2& 	1	 &    & \\
  		\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n} \\
  		-x_n& 	1 &  & \\
  	\end{array} \right].
  \end{align}
Note that $\det G =\det \tilde{ G}$.
Thus
\begin{align}
	\det (	G (\boldsymbol{\alpha} \boldsymbol{\alpha}^T +\tilde{ U}	) G)\nonumber
=	&\	\det  \tilde{G }^{-1}L\tilde{G }^{-1}	)   \nonumber\\
=&\	\det\left[\begin{array}{cc|cccc}
		0 & 1 & x_1&x_2& \cdots & x_n \\
		-1& y  &1  &1 & \cdots &  1\nonumber \\
		\hline
		-x_1& 1  & &\nonumber \\
		-x_2& 	1	 &    & \nonumber \\
		\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}\nonumber \\
		-x_n& 	1 &  &\nonumber \\
	\end{array} \right]   \nonumber  \\
&\ +\det\left[\begin{array}{cc|cccc}
	1 & 1 & x_1&x_2& \cdots & x_n   \nonumber\\
	0& y  &1  &1 & \cdots &  1    \nonumber\\
	\hline
	0& 1  & &   \nonumber \\
0& 	1	 &    &      \nonumber\\
	\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}      \nonumber \\
	0& 	1 &  &      \nonumber\\
\end{array} \right]        \nonumber\\
\nonumber
\end{align}
and hence
\begin{equation}\label{alph+u}\begin{aligned}
&\	\det (	G(\boldsymbol{\alpha} \boldsymbol{\alpha}^T +\tilde{ U}	) G) -\det\left[\begin{array}{c|cccc}
	 y & 1& \cdots & 1 \\
	 	\hline
		1	 &    & \\
		\vdots & &\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n} \\
		1 &  & \\
\end{array} \right]
\\=&\
 -\det\left[\begin{array}{cc|cccc}
	0 & 1 & x_1&x_2& \cdots & x_n        \\
	1& y  &1  &1 & \cdots &  1       \\
	\hline
	x_1& 1  & &      \\
	x_2& 	1	 &    &        \\
	\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}        \\
	x_n& 	1 &  &        \\
\end{array} \right].         \\
\end{aligned}\end{equation}
Combining (\ref{u}) with (\ref{alph+u}),  we obtain
\begin{equation}\label{GGG}\begin{aligned}
	&\	\det (	G (\boldsymbol{\alpha}  \boldsymbol{\alpha}^T +\tilde{ U}	) G) -\det (G  \tilde{ U}	 G)\\
	=&\	-\det\left[\begin{array}{cc|cccc}
		0 & 1 & x_1&x_2& \cdots & x_n\\
		1& y  &1  &1 & \cdots &  1 \\
		\hline
		x_1& 1  & & \\
		x_2& 	1	 &    & \\
		\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n} \\
		x_n& 	1 &  &
	\end{array} \right].
\end{aligned}\end{equation}
Exchanging the first two rows and  the first two columns, the last determinant turns out to be
\begin{align}
&\ \det\left[\begin{array}{cc|cccc}
	0 & 1 & x_1&x_2& \cdots & x_n \nonumber\\
	1& y  &1  &1 & \cdots &  1\nonumber \\
	\hline
	x_1& 1  & &\nonumber \\
	x_2& 	1	 &    & \nonumber\\
	\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n} \nonumber\\
	x_n& 	1 &  &\nonumber \\
\end{array} \right]	
\\
=&\	\det\left[\begin{array}{cc|cccc}
		y & 1&1  &1 & \cdots &  1 \nonumber \\
		1& 0   & x_1&x_2& \cdots & x_n \nonumber \\
		\hline
	 1  &	x_1& &\nonumber \\
		1	 & 	x_2&    &\nonumber \\
		\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}\nonumber \\
			1 & x_n&  &\nonumber \\
	\end{array} \right]	
\\
=&\	\det\left[\begin{array}{cc|cccc}
		1 & 1&1  &1 & \cdots &  1 \nonumber \\
		1& 0   & x_1&x_2& \cdots & x_n \nonumber \\
		\hline
		1  &	x_1& &\nonumber \\
		1	 & 	x_2&    &\nonumber \\
		\vdots & 	\vdots && &\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}\nonumber \\
		1 & x_n&  & \\
	\end{array} \right]	
\\&\ +	\det\left[\begin{array}{cc|cccc}
		y -1& 1&1  &1 & \cdots &  1 \nonumber \\
		0& 0   & x_1&x_2& \cdots & x_n\nonumber  \\
		\hline
		  &	x_1& & \\
		0	 & 	x_2&    & \\
		\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n} \nonumber\\
		0& x_n&  &\nonumber
	\end{array} \right].	
\end{align}
Thus, in view of \eqref{GGG}, we have
\begin{align*}
	&\	\det (G  \tilde{ U}	 G)-\det (	G  (\boldsymbol{\alpha}  \boldsymbol{\alpha}^T +\tilde{ U}	) G) \\
	=&\ \det\left[\begin{array}{cc|cccc}
		1 & 1&1  &1 & \cdots &  1     \nonumber  \\
		1& 0   & x_1&x_2& \cdots & x_n \nonumber \\
		\hline
		1  &	x_1& &\nonumber \\
		1	 & 	x_2&    &\nonumber \\
		\vdots & 	\vdots & &&\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}           \nonumber \\
		1 & x_n&  & \nonumber\\
	\end{array} \right]	 \nonumber
\\&\
	+(y-1) \det \left[ \begin{array}{c|cccc}
		0& x_1 & x_2  & \cdots & x_n\nonumber \\
		\hline
		x_1 & & & \nonumber\\
		x_2	 & &   &\nonumber \\
		\vdots & & &\bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{1\le i,j\le n}\nonumber \\
		x_n& & &\nonumber \\
	\end{array} \right].
\end{align*}

Let $x_{-1}=1$ and $x_{0}=0$. Then
the last two determinants are
 \begin{equation*}
 	\det \bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{-1\le i,j\le n}
 \ \t{and}\
  \det \bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{0\le i,j\le n}
\end{equation*}
respectively.
Thus
\begin{align}\label{alphaU-U}
	 &	\det (	G (\boldsymbol{\alpha}  \boldsymbol{\alpha}^T +\tilde{ U}	)G) -\det (G  \tilde{ U} G)  \nonumber\\
=	&- \det \bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{-1\le i,j\le n} -(y-1)\det \bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{0\le i,j\le n} .
\end{align}
To evaluate the right side of (\ref{alphaU-U}), we define
\begin{align}
	I =\det \bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{-1\le i,j\le n}
\ \t{and}\
	J =\det \bigg[\frac{x_i+x_j}{1+x_ix_j}\bigg]_{0\le i,j\le n} .
\end{align}
It is easy to see that
\begin{align}\label{I+J}
	\det (	G(\boldsymbol{\alpha}  \boldsymbol{\alpha}^T +\tilde{ U}	)G) -\det (G  \tilde{ U} G)= - I-(y-1)J.
\end{align}
By Lemma \ref{uv},
\begin{align*}
I=	(-1)^{\frac{p+3}{4}} \prod_{i=1}^{n}(  1-x_i )^2
	\cdot\prod_{1\le i<j\le n}^{}  (x_i-x_j) \prod_{i=1}^{n}  \prod_{j=1}^{n}(1+x_ix_j)^{-1} \prod_{i=1}^{n}x_i^2
\end{align*}
and
\begin{align}
 J	=&\ (-1)^{\f{p-1}4}\cdot\frac{1}{2}\(\prod_{i=1}^{n}(  1+x_i )^2- \prod_{i=1}^{n}(  1-x_i )^2\)   \nonumber\\
	&\ \cdot\prod_{1\le i<j\le n}^{}(x_i-x_j)\cdot\prod_{i=1}^{n}  \prod_{j=1}^{n}(1+x_ix_j)^{-1}\prod_{i=1}^{n}x_i^2.\nonumber
\end{align}
Set
\begin{equation*}
	f_1= \prod_{1\le i<j\le n}^{}\(   \(\frac{j}{p}\)\zeta^{j}- \(\frac{i}{p}\)\zeta^{i} \)
	= \prod_{1\le i<j\le n}^{}(  x_j^{-1}- x_i^{-1} )
\end{equation*}
and
\begin{equation*}
	f_2= \prod_{1\le i<j\le n}^{}\( 1+  \(\frac{j}{p}\)\zeta^{j}\(\frac{i}{p}\)\zeta^{i} \)	= \prod_{1\le i<j\le n}^{}( 1+ x_j^{-1} x_i^{-1} ).
\end{equation*}
Then
\begin{align}
	\prod_{1\le i<j\le n}^{} (x_i-x_j) \cdot\prod_{i=1}^{n}  \prod_{j=1}^{n}(1+x_ix_j)^{-1}\cdot  \prod_{i=1}^{n}x_i^2= f_1^2\  f_2^{-2} \prod_{i=1}^{n}\l((1+ x_i^{2})^{-1}x_i^2\r).
	\nonumber
\end{align}
By \cite[Corollary 2 (4.6)]{V1}, we have
\begin{align}
	\prod_{i=1}^{n}(1+ x_i^{2})^{-1}\cdot  \prod_{i=1}^{n}x_i^2=\zeta^{-n(n+1)/2} \(\frac{2}{p}\).  \nonumber
\end{align}
So
\begin{align*}
	\prod_{1\le i<j\le n}^{} (x_i-x_j) \cdot\prod_{i=1}^{n}  \prod_{j=1}^{n}(1+x_ix_j)^{-1}\cdot  \prod_{i=1}^{n}x_i^2= f_1^2 f_2^{-2} \zeta^{-n(n+1)/2} \(\frac{2}{p}\).	\nonumber\\
\end{align*}
Then, in light of Lemma \ref{a+-b}, we have
$$
I	=  ( a_p' \sqrt{p}-b_p' p)  f_1^2  f_2^{-2} \zeta^{-n(n+1)} \(\frac{2}{p}\)$$
and
$$
	J	=  a_p' \sqrt{p}\, f_1^2  f_2^{-2} \zeta^{-n(n+1)} \(\frac{2}{p}\).$$
 Combining the above results with (\ref{I+J}),  and noting that $y= (\frac{2}{p})\sqrt{p}\,x$
 and  $	(\det G)^2= \zeta^{n(n+1)} $, we obtain
\begin{align*}
&\ \det (\boldsymbol{\alpha}  \boldsymbol{\alpha}^T +\tilde{ U}	)- \det \tilde{ U}
\\=&\  \zeta^{-n(n+1)}(-I-(y-1)J)\\
=&\  \(\(\frac{2}{p}\)b_p'-a_p'x \) p\zeta^{-2n(n+1)}   f_1^2  f_2^{-2}.
\end{align*}
This, together with  (\ref{Rx})  and (\ref{alphaUalpha}), yields that
\begin{align*}
	 \det R(x)
	 =&\ \(\(\frac{2}{p}\)b_p'-a_p'x \) \zeta^{\frac{(p-1)^2}{8}}  p^{\frac{p+3}{4}}
 (\det V)^2  (\det D)^2 f_1^2  f_2^{-2}      \\
	 =&\ \l(\frac{2}{p}\r)b_p'-a_p'x \quad \ (\t{by \cite[(4.10)]{V1}}).
\end{align*}
This completes our proof of \eqref{simple} for $p\eq1\pmod4$. \qed

\medskip
\noindent{\bf Statements and Declarations}. There are no competing interests. This original paper contains no data, and  it has not been submitted elsewhere.

\end{document}